4. Your company has 50 employees who receive 28 vacation days per year. They work 8 hour days and make $20 per hour. Vacation days not taken are carried over to the next year. Here are the days earned and taken in 2019: 

Employees Days earned and not taken Carryover days 10 0 0 20 7 140 20 07 100 = 50 = 240 

a. Record vacation taken in 2019 

b. Estimate and record the liability for compensated absences 

c. Record vacation actually taken if 100,000 was the actual cost
Answer #1

a) No. of vacations available per employee during the year = 28

Total no. of employees = 50

Total no. of vacation days available = 28×50= 1400 days

Vacation days carryover = 240 days

Total vacation taken in 2019 = 1400-240= 1160 days, which can be further explained as below:

For 10 employees, vacation days taken = 10×28= 280 days

For 20 employees, vacations days taken = 20×21= 420 days

For next 20 employees, vacation days taken = 20×23=460 days